y = 2(x - (-4))2 + (-7)
Which choice converts this standard parabolic equation to its simplified form?
A. y = 2x2 + 12x + 12
B. y = 2x2 + 16x + 25
C. y = 4x2 + 4x + 8
D. x = 2y2 + 8y + 12

Answer by lwsshak3(6505)

y = 2(x - (-4))2 + (-7)
y=2(x+4)^2-7
y=2(x^2+8x+16)-7
y=2x^2+16x+32-7
y=2x^2+16x+25
B is the right answer